Raj is 11 years old and his uncle is 59 years . How many years ago was raj's uncle 7 times as old as raj.
Answer by swincher4391(1107)

Let x be how many years ago Raj's uncle was 7 times older than raj.
Then we have 7*(11-x) = (59-x) (since we went back in time x years)
77 - 7x = 59 - x
18 = 6x
x = 3
So, 3 years ago. Let's check:
3 years ago Raj was 8 and his uncle was 56. 8x7=56. Correct.